1/1

McGregor of Glen Endrick,

The Bonny Banks of Loch Lomond

By yon bonnie banks and by yon bonnie braes Where the sun shines bright on Loch Lomond Me and my true love were ever wont to gae On the bonny, bonny banks of Loch Lomond

Chorus Ye’ll tak’ the high road and I’ll tak the low road And I’ll be in Scotland afore ye But me and my true love will never meet again On the bonny, bonny banks of Loch Lomond

’Twas there that we parted in yon shady glen On the steep, steep side of Ben Lomond Where in the purple hue the hieland hills we view And the moon coming out in the gloaming

The wee birdies sing and the wild flowers spring And in sunshine the waters are sleeping But the broken heart it kens nae second spring again And the waefu’ may cease frae their Th

Verkinto de tiu ĉi Scottish poemo estas McGregor of Glen Endrick.

Arg-689-1390 (2012-12-03 10:23:46)

Pri la tekstso vidu: http: // www. darachweb. net/ SongLyrics/ LochLomond. html . Pri la muziknotoj vidu:http: // mudcat. org/ @displaysong. cfm? SongID= 3667 . La melodio estas aŭdebla en http: // www. youtube. com/ watch? v= sAWg02PZelQ . Vidu ankaŭ la version verkitan de la poeto Andrew Lang (1844 - 1912): https: // tspace. library. utoronto. ca/ html/ 1807/ 4350/ poem2530. html .

http://www.poezio.net/version?poem-id=689&version-id=,1390